### v1

# Spin Squeezing Enhanced Quantum Magnetometry with Nitrogen-Vacancy Center Qutrits

Abstract
We explore the utility of quantum spin squeezing in quantum magnetometry, focusing on three-level (qutrit) Nitrogen-Vacancy (NV) centers within diamond, utilizing a standard Ramsey interferometry pulse protocol. Our investigation incorporates the effects of dephasing and relaxation on NV centers’ dynamics during Ramsey measurements, modeled via the Lindblad quantum master equation. We conduct a comparative analysis between the metrological capabilities of a single NV center and a pair of NV centers, considering Quantum Fisher Information both with and without spin squeezing. The quantum correlations between NV centers are assessed through the evaluation of the Kitagawa-Ueda spin squeezing parameter within a two-level manifold. Additionally, parallel calculations are conducted using a two-level model (qubit) for NV centers. Our findings reveal that leveraging qutrits and spin squeezing yields enhanced magnetometric precision, albeit constrained by dephasing effects. Nevertheless, even in the absence of dynamical decoupling methods to mitigate environmental noise, strategic timing of squeezing and free evolution can sustain the advantages of qutrit-based magnetometry.
I Introduction
The exploration of precisely controlled quantum systems as highly sensitive nanoscale detectors holds significant promise in advancing our comprehension of intricate processes within biological and condensed-matter systems, operating at molecular and atomic scales [1, 2]. The demanding criteria for both high sensitivity and spatial resolution have prompted suggestions to employ spin-based quantum systems as nanoscale magnetometers [3] or to utilize imaging through the detection of sample-induced decoherence [4]. A particularly appealing physical platform for realizing these concepts is the Nitrogen-Vacancy (NV) center in diamond. This choice is motivated by the center’s prolonged coherence times at room temperature and its convenient optical readout of the spin state [5, 6, 7, 8, 9, 10].
Given the considerable potential of the NV center, researchers have conducted studies to enhance the sensing capabilities of this probe. Specific pulse sequences, such as Carr-Purcell-Meiboom-Gill (CPMG) and Dynamic Decoupling (DD) [11, 12], have been used to prolong coherent time. Additionally, researchers have investigated the use of entangled NV centers [13, 14] and, finally, spin squeezing [15]. Spin squeezing, as introduced by Kitagawa and Ueda [16], involves the quantum redistribution of uncertainties along two orthogonal spin directions. This operation holds significance in quantum sensing and metrology, providing a means to enhance measurement precision beyond the standard quantum limit in experiments [17, 18, 19]. Consequently, spin squeezing emerges as a valuable resource in applications of quantum technology. In the context of spin squeezing generation for NV centers, ensembles of qubits are typically considered. However, this approach has limitations as the qubit model is an approximation that necessitates a bias field. Here, we explore the potential of using qutrits for NV centers. Pulsed dynamical generation of spin squeezing has been explored in the literature for spin-1 systems [20, 21] and other methods, including generation via nonadiabatic control [22]. In the context of NV centers, Ramsey interferometry is recognized as a straightforward and efficient protocol for measuring magnetic fields. It has been extensively studied in qubit systems [23, 24, 25, 26] and has also been discussed for qutrit systems [27, 28].
In this study, we explore the potential enhancements in quantum metrological performance achievable by transitioning from two-level (qubits) to three-level (qutrits) spin systems and incorporating spin squeezing within typical quantum magnetometry settings, specifically NV center Ramsey interferometry. We adopt a foundational approach, comparing qubit versus qutrit Ramsey quantum magnetometry scenarios, both with and without spin squeezing, in terms of Quantum Fisher Information (QFI). Considering the influence of quantum dephasing and relaxation during the spin system’s evolution, our findings suggest that qutrits, coupled with the additional step of spin squeezing, can outperform qubits in terms of QFI, provided that the spin squeezing magnitude, the timing of squeezing and free evolution are optimally determined.
The structure of the paper is as follows: Section II provides an overview of the employed model, elucidating the Hamiltonian and dynamics. In Section III, we introduce QFI and spin squeezing parameter. Section IV presents our results, and finally, Section V summarizes our conclusions.
II Model
In a scenario where there exists an unknown, potentially weak magnetic field within a defined region, our objective is to measure this magnetic field with the utmost precision possible. To accomplish this, we utilize one or two NV centers as our probes. The NV center, characterized as an optically active color defect, results from a substitutional nitrogen impurity and a neighboring carbon vacancy in the diamond lattice. Comprising six electrons from nitrogen and surrounding carbon atoms, the negatively charged NV center is described as a spin-1 system. We opt to explore the full capabilities of the NV center by modeling it as a qutrit system. This choice is motivated by the potential advantages that qutrits offer in metrology when compared to qubits [29, 30]. Its ground state is a spin triplet (${}^{3}A_{2}$) denoted as |$\ket{Sm_{S}}$ with $S=1$ and $m_{S}=0,\pm 1$. The NV center Hamiltonian captures these properties, while the excited-state triplet (${}^{3}E$) is acknowledged at a higher energy level and is not explored in this context [31, 32].
Initially, our focus is on a single NV center. We then extend our consideration to two NV centers, exploring Ramsey interferometry with and without the application of spin squeezing as illustrated in Fig. 1.
Figure 1: Illustration of Ramsey-type sequence with and without introduction of spin squeezing. Note that the times mentioned in the figure correspond to those considered in our study.
II.1 Free Hamiltonian and Initial states
Neglecting interactions with nuclear spins or spin-strain effects [33, 34], the Hamiltonian governing $N$ NV centers is expressed as follows,
$\begin{split}\hat{H}_{0}&=\sum_{i=1}^{\text{N}}\left[D\hat{S}_{zi}^{2}+g_{S}\vec{B}\cdot\vec{\hat{S}}_{i}\right].\end{split}$ (1)
We take $\hbar=1$ here and in the rest of the paper. $D=2.87$ GHz is the temperature-dependent zero-field splitting, and $g_{S}=1.80$ MHz/G is the gyromagnetic ratio for the electron spin. The magnetic field taken parallel to the NV center axis ($z$) is denoted as $\vec{B}=(0,0,B_{z})$. The spin-1 operator $\vec{\hat{S}}=(\hat{S}_{x},\hat{S}_{y},\hat{S}_{z})$ is defined as
$\hat{S}_{x}=\frac{\hbar}{\sqrt{2}}\begin{bmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{bmatrix},$ (2)
$\hat{S}_{y}=\frac{\hbar}{\sqrt{2}}\begin{bmatrix}0&-i&0\\ i&0&-i\\ 0&i&0\end{bmatrix},$ (3)
$\hat{S}_{z}=\hbar\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&-1\end{bmatrix}.$ (4)
$\hat{S}_{i}$ is the spin matrix for the $i^{th}$ NV center. Due to the Zeeman effect in the presence of a magnetic field, the electron spin ($S=1$) is split into three states: $m_{s}=0$, $1$, $-1$, denoted as |$\ket{0}$, |$\ket{1}$ and |$\ket{-1}$. The initial state of the NV center is taken in the excited state |$\ket{-1}$,
|$\ket{\psi_{0}}=\bigotimes_{i=1}^{N}\ket{-1}.$ (5)
Then, each NV center undergoes a $\pi/2$ pulse, achieved by applying a rotation operator,
$\mathrm{e}^{-i\frac{\pi}{2}S_{x}}.$
This pulse is experimentally generated by microwave drives [35].
When incorporating spin squeezing (details provided in the following subsections), the initial state, |$\ket{\psi_{0}}$, is first driven through spin squeezing, and then subjected to the $\pi/2$ pulse, as illustrated in Fig. 1.
II.2 Spin Squeezing Generation
To describe the intricate dynamics and capture the squeezing effects in the Hamiltonian for two NV centers, a squeezing term is introduced [36],
$\hat{H}_{\mathrm{S}}=c_{1}\left(\sum_{i=1}^{N}\hat{S}_{xi}\right)^{2},$ (6)
where $c_{1}$ is the squeezing constant. The Hamiltonian, as depicted in Eq. (6), is referred as a one-axis twisting Hamiltonian. The choice of $x$-axis twisting in the context of spin squeezing is strategic, as it allows for the manipulation of the noise cone, effectively narrowing the error ellipse of the spin. This, combined with the optimal ground state, enhances the precision of the magnetic field measurement by interplaying with the $z$-axis rotation induced by the magnetic field [37].
The generation of this type of Hamiltonian in real setup has been discussed in the literature. It can be realized through phonon mediation, as evidenced in [38] or produced in optical cavities [39]. In our specific protocol, we need to deactivate spin squeezing after a specific duration. A dynamic control method for NV center spin squeezing, as discussed in [40], could potentially address the technical challenge.
Alternative approaches to generating spin squeezing have been proposed. One example involves using microwave fields in silicon-vacancy centers, which are similar to NV centers [41]. Additionally, a universal method, introduced in [21], is also applicable to NV centers. The method presents a generic scheme for creating spin squeezing in any non-squeezing systems through the use of periodic rotation pulses.
Figure 2: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G. The curves represents the free evolution as follows: the solid blue curve corresponds to a single NV center, the dashed-dotted red curve corresponds to two NV centers, the pointed cyan curve represents two NV centers subjected to spin squeezing ($c_{1}=2.5$ GHz, $T_{\mathrm{sq}}=3$ ns), and the dashed green curve represents two NV centers subjected to spin squeezing ($c_{1}=0.25$ GHz, $T_{\mathrm{sq}}=30$ ns). The different scenarios are illustrated as: (a) without including any noise terms, (b) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=0.1\gamma_{t}$, and (c) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=\gamma_{t}$.
II.3 Dynamics
We describe the open quantum system dynamics of the NV center by the following Lindblad master equation [42],
$\dot{\rho}(t)=-i[\hat{H},\rho(t)]+\mathcal{D}_{t}(\rho(t))+\mathcal{D}_{d}(\rho(t)),$ (7)
where the unitary contribution to the dynamics, $\hat{H}$, is provided as follows,
  • during free evolution by $\hat{H}=\hat{H}_{0}$,
  • during spin squeezing by $\hat{H}=\hat{H}_{0}+\hat{H}_{\mathrm{S}}$.
The NV center, while a robust quantum system, is inevitably influenced by the fluctuating states of nearby nuclear spins. These nuclear spins introduce a dynamic and unpredictable backdrop, contributing to the environmental noise that impacts the precision of measurements performed on the NV center. We assume very low temperatures relative to the NV center transition frequencies. Dephasing and thermalization are incorporated through the dissipative rate. They are respectively given as follows,
$\mathcal{D}_{d}(\rho)$ $=\sum^{N}_{i=1}\gamma_{\mathrm{d}}\,[\hat{S}_{zi}\rho(t)\hat{S}_{zi}-\rho(t)];$ (8)
$\mathcal{D}_{t}(\rho)$ $=\sum^{N}_{i=1}\gamma_{t}\,\left[\hat{S}_{i}^{-}\rho(t)\hat{S}_{i}^{+}-\frac{1}{2}\left\{\hat{S}_{i}^{+}\hat{S}_{i}^{-},\rho(t)\right\}\right],$ (9)
where $\hat{S}_{i}^{-}$ is the lowering operator of the $i^{\text{th}}$ NV center, and,
$\hat{S}^{-}$ $=\frac{1}{2}(\hat{S}_{x}-i\hat{S}_{y}).$ (10)
[…]
III Quantum Measures
III.1 Quantum Fisher Information
In the context of quantum parameter estimation, particularly for magnetic field sensing, fundamental results in quantum metrology become imperative. Let us consider a scenario where a quantum state encodes a parameter $\theta$, and the goal is to estimate this parameter with optimal precision [44, 45]. For a mixed state described by the density operator, $\hat{\rho}$, the QFI can be defined, using the Symmetric Logarithmic Derivative (SLD) operator ($\hat{L}_{\theta}$), as
$F_{Q}(\theta)=\text{Tr}[\hat{\rho}\hat{L}_{\theta}^{2}].$ (11)
The SLD operator is implicitly determined by the equation,
$\partial_{\theta}\hat{\rho}=\frac{\hat{L}_{\theta}\hat{\rho}+\hat{\rho}\hat{L}_{\theta}}{2}.$ (12)
Expressing $\hat{L}_{\theta}$ in the eigenbasis of $\hat{\rho}$, the QFI takes the form,
$F_{Q}(\theta)=2\sum_{k,l}\frac{\left|\langle k|\partial_{\theta}\hat{\rho}|l\rangle\right|^{2}}{\lambda_{k}+\lambda_{l}}.$ (13)
Here, |$\ket{k}$ and |$\ket{l}$ are the eigenvectors and $\lambda_{k}$ and $\lambda_{l}$ are the eigenvalues of the density operator $\hat{\rho}$.
In the specific application of estimating a magnetic field, $\theta$, in our case, is the strength of the magnetic field, and the QFI provides a quantitative measure of the ultimate precision achievable in such quantum magnetic field sensing experiments. This theoretical foundation is critical for understanding the limits and potential advancements in quantum-enhanced magnetic field measurements. To quantify this precision, we employ the QuanEstimation toolbox in Python [46] to calculate the QFI.
III.2 Spin Squeezing Parameter
The concept of spin squeezing is established by defining it as the condition where the variance of a spin component normal to the mean spin is smaller than the standard quantum limit. To quantify the extent of spin squeezing, we employ the spin squeezing parameter, as introduced by Kitagawa and Ueda [16]. Kitagawa-Ueda spin squeezing brings quantum entanglement perspective to the usual spin squeezing which is the reduction of spin noise in a certain direction at the cost of more noise in other directions. This criterion is given by
$\xi^{2}=\frac{4(\Delta\hat{J}^{2}_{n_{\perp}})_{\text{min}}}{N},$ (14)
where $n_{\perp}$ is an axis perpendicular to the mean-spin direction $\hat{n}$, $N$ is the number of spins (in our case, $N=2$), and $\hat{J}$ represents the collective spin operator. Kitagawa-Ueda and other quantum correlations based criteria of spin squeezing [47] are commonly defined for SU(2) or spin-1/2 atomic ensembles. Spin squeezing in SU(3) or spin-1 systems can be discussed in their SU(2) subgroups, such as isospin squeezing [48]. We follow a similar approach to characterize quantum correlations and spin squeezing within the manifold of {|$\{\ket{0},\ket{-1}\}$. In general, spin squeezing can be optimized by rotating the coordinate frame and choosing the critical rotation angles. For simplicity, we only consider a fixed reference frame. Based on this manifold, the spin matrices $\hat{\sigma}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ are written as
[…]
and the collective spin $\hat{J}=(\hat{J}_{x},\hat{J}_{y},\hat{J}_{z})$ is defined as
$\hat{J}_{x}$ $=\frac{1}{2}(\hat{\sigma}_{x}\otimes I_{3}+I_{3}\otimes\hat{\sigma}_{x}),$ (18)
$\hat{J}_{y}$ $=\frac{1}{2}(\hat{\sigma}_{y}\otimes I_{3}+I_{3}\otimes\hat{\sigma}_{y}),$ (19)
$\hat{J}_{z}$ $=\frac{1}{2}(\hat{\sigma}_{z}\otimes I_{3}+I_{3}\otimes\hat{\sigma}_{z}),$ (20)
with $I_{3}$ being the three dimensional unit matrix. To calculate spin squeezing, we calculate the mean spin squeezing direction $\hat{n}$ for the state and then take the minimal variance in the perpendicular direction of $\hat{n}$,
$\hat{n}=\frac{(\langle\hat{J}_{x}\rangle,\langle\hat{J}_{y}\rangle,\langle\hat{J}_{z}\rangle)}{|\langle\hat{J}\rangle|}.$ (21)
The presence of spin squeezing is characterized by $\xi^{2}<1$. Spin squeezing is closely tied to the creation of entanglement [37, 49, 50]. Additionally, a previous study suggests that initial squeezing can aid in entanglement [51]. Our results are consistent with this statement (Appendix A).
Figure 3: The curve in solid blue represents the squeezing parameter, $\xi^{2}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
IV Results
Initially, we investigate the typical Ramsey sequence for one and two NV centers, both starting in the initial state (Eq. 5). Subsequently, after a $\pi/2$ pulse, we allow the system to undergo evolution through the free Hamiltonian $\hat{H}_{0}$ for a duration of $T_{\mathrm{fr}}=20$ $\mu$s, $T_{\mathrm{fr}}$ being the free evolution time. We choose this time taking into account the dissipative time that falls within the order of tenth of microseconds. To further enhance precision, starting in the initial state (Eq. 5), we consider introducing spin-squeezing before the $\pi/2$ pulse and the free evolution. The two scenarios are illustrated in Fig. 1.
IV.1 Analysis with spin-1 formalism
[…]
Fig. 2 compares two NV centers under different conditions: one without spin squeezing and the other with spin squeezing applied at varying strengths, specifically with squeezing constants of $c_{1}=0.25$ GHz and $c_{1}=2.5$ GHz. In the absence of noise, the QFI is larger when spin squeezing is applied, and it increases further with a higher squeezing constant, $c_{1}$ (Fig. 2a). When environmental noise is introduced with $\gamma_{\text{d}}=0.1\gamma_{\text{t}}$, spin squeezing remains beneficial, particularly at a lower squeezing constant ( $c_{1}=0.25$ GHz) (Fig. 2b). Applying moderate squeezing, sufficiently large to achieve its effects, shows advantages by being more robust against the impact of environmental noise. However, within the first 4 $\mu$s, a higher squeezing constant provides more benefit. As the dephasing rate increases to $\gamma_{\text{d}}=\gamma_{\text{t}}$, the advantage of squeezing is lost (Fig. 2c).
The effect of spin squeezing on the NV centers is further illustrated in Fig. 3. The squeezing parameter, initially around 0.5 (below 1) at the beginning of free evolution, increases due to dephasing and thermalization during this period. Appendix A demonstrates that spin squeezing also leads to entanglement. The duration of the application of spin squeezing, denoted by $T_{\mathrm{sq}}$, corresponds to the time when the squeezing parameter reaches its minimum value.
The initial state is oriented along the $z$-axis. When spin squeezing is applied, it reduces the variance in a specific direction, creating an ellipsoid shape that remains aligned with the $z$-axis. A $\pi/2$ pulse then shifts the spin into the $x$-$y$ plane while preserving this variance ellipsoid. The effect of the magnetic field $B$ is to precess the spin around the $z$-axis, and the $x$-axis twisting deforms the noise cone, effectively narrowing the spin’s error ellipse. This sequence is illustrated in Fig. 4.
Figure 4: Illustration of the mean spin orientation and the variance of state preparation before the free evolution of the NV centers. At time $t_{1}$, the state is in its initial configuration (Eq. (5)) with homogeneous variance. At time $t_{2}$, the state reaches the end of the spin squeezing period, remaining aligned in the $z$-direction but with noise squeezed into an ellipse. Finally, at $t_{3}$, a $\pi/2$ pulse is applied, shifting the state into the $x$-$y$ plane.
Experimentally, we typically measure populations or ($\hat{S}_{z}$) values. However, it is important to note that while the QFI is calculated for the optimal Positive Operator-Valued Measure (POVM), specifically the eigenprojectors of the Symmetric Logarithmic Derivative (SLD), measuring $\hat{S}_{z}$ could potentially be suboptimal.
While spin squeezing enhance magnetic sensing capabilities, they are also highly sensitive to environmental noise, particularly dephasing. Analyzing the behavior of the QFI curves under such noise conditions is crucial. By identifying the time at which the QFI reaches its maximum, we can optimize experimental design and ensure measurements are performed at the peak sensitivity.
Figure 5: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G. The curves represents the free evolution as follows: the solid blue curve corresponds to a single NV center, the dashed-dotted red curve corresponds to two NV centers, the pointed cyan curve represents two NV centers subjected to spin squeezing ($c_{1}=2.5$ GHz, $T_{\mathrm{sq}}=3$ ns), and the dashed green curve represents two NV centers subjected to spin squeezing ($c_{1}=0.25$ GHz, $T_{\mathrm{sq}}=30$ ns). The different scenarios are illustrated as: (a) without including any noise terms, (b) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=0.1\gamma_{t}$, and (c) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=\gamma_{t}$. The NV centers system are described by spin-1/2 formalism.
IV.2 Analysis with spin-1/2 formalism
In this section, we examine the outcomes within the identical parameter scope for the spin-1/2 framework. This is achieved by concentrating on the dynamics of the states |$\ket{0}$ and |$\ket{-1}$. The Hamiltonian, in this context, is characterized as follows,
$\hat{H}_{0}=\sum_{i=1}^{N}\frac{\omega}{2}\hat{\sigma}_{zi},$ (22)
where $\omega=D-g_{S}B_{z}$ and $\hat{\sigma}_{i}=(\hat{\sigma}_{xi},\hat{\sigma}_{yi},\hat{\sigma}_{zi})$ are the Pauli matrices for the $i^{\text{th}}$ NV center. The parameters and open quantum system dynamics are kept same as in the spin-1 model. Eq. (8) and Eq. (9) are modified respectively to
$\mathcal{D}_{d}(\rho)$ $=\sum^{N}_{i=1}\gamma_{\mathrm{d}}\,[\hat{\sigma}_{zi}\rho(t)\hat{\sigma}_{zi}-\rho(t)];$ (23)
$\mathcal{D}_{t}(\rho)$ $=\sum^{N}_{i=1}\gamma_{t}\,[\hat{\sigma}_{i}^{-}\rho(t)\hat{\sigma}_{i}^{+}-\frac{1}{2}\left\{\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-},\rho(t)\right\}],$ (24)
where $\hat{\sigma}_{i}^{\pm}$ are the Pauli spin ladder operators for the $i^{\text{th}}$ NV center. The spin squeezing Hamiltonian in Eq. (6) is modified to
$\hat{H}_{\mathrm{S}}=c_{1}\left(\sum^{N}_{i=1}\hat{\sigma}_{xi}\right)^{2}.$ (25)
The initial state of the NV center is taken in the excited state |$\ket{-1}$. Then each NV center undergoes an $\pi/2$ pulse, achieved by applying a rotation operator,
$\mathrm{e}^{-i\frac{\pi}{2}\hat{\sigma}_{x}}.$
As in the spin-1 formalism, if spin squeezing is added, it is applied before the $\pi/2$ pulse. We then proceed to analyze the QFI under both scenarios, considering the presence and absence of noise.
In Fig. 5a, we observe that without considering any noise channel, spin squeezing yields a higher QFI. However, in the presence of noise (Fig. 5b and Fig. 5c), the advantages of spin squeezing are lost even for small dephasing (Fig. 5b), and Ramsey sequence without spin squeezing is more beneficial. In the case of the qutrit, we observe advantages even in the presence of dephasing (Fig. 2b). Therefore, we can see the benefits of using qutrit sensors over qubit sensors.
V Conclusion
We have examined the potential of Nitrogen-Vacancy (NV) centers as quantum sensors for magnetic fields, exploring various scenarios involving single and paired NV centers using both qubit and qutrit models, and employing spin squeezing within Ramsey or Ramsey-type pulse sequence protocols. Throughout our investigation, we have analyzed the dynamics of Quantum Fisher Information (QFI) across these scenarios, utilizing the Lindblad master equation to model the spin system’s evolution under dephasing and thermalizing noise.
Incorporating a spin squeezing step in magnetometric measurement and utilizing qutrit NV centers have demonstrated a precision increase relative to standard Ramsey magnetometry with qubits and without spin squeezing. However, the advantage of qutrit spin squeezing is notably sensitive to the dephasing environment. Our numerical analysis showed that, as dephasing increases, the advantages become marginal compared to standard methods. We also observed that moderate squeezing results in higher QFI compared to strong squeezing, which is only beneficial for a brief period of free evolution.
Additionally, we have observed that spin squeezing is accompanied by the generation of entanglement, aligning with the general expectation of quantum entanglement as a metrological resource. However, further investigation is needed to determine whether it is spin noise distribution or entanglement alone that causes the precision limit enhancement. Our findings contribute to the broader exploration of quantum technologies beyond qubits, where qutrits and higher-dimensional qubits (qudits) are increasingly considered, particularly in the context of quantum computing [52, 53]. An intriguing question persists regarding the scalability of these advantages, especially regarding the utilization of NV center qutrit ensembles or the investigation of higher-level systems (qudits). These avenues pose exciting challenges for future studies, promising further insights.
Figure 6: The curve in solid blue represents the logarithmic negativity, $E_{N}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
Appendix A Entanglement Measure
For higher-dimensional systems than qubits, some entanglement measures exist. The negativity, denoted as $N(\rho_{AB})$, quantifies entanglement in bipartite quantum systems $\rho_{AB}$ [54]. It is defined as the absolute sum of the negative eigenvalues of the partially transposed density matrix $\rho_{A}^{T}$,
$N(\rho_{AB})=\sum_{i}|\lambda_{i}|,$ (26)
where $\lambda_{i}$ are the negative eigenvalues of the density matrix $\rho_{A}^{T}$. The logarithmic negativity, denoted as $E_{N}(\rho_{AB})$, is related to the negativity and serves as a good indicator of the degree of entanglement in our system. It is defined as
$E_{N}(\rho_{AB})=\log_{2}(2N(\rho_{AB})+1).$ (27)
In Fig. 6, we observe that the application of spin squeezing to the two NV centers system creates an entanglement between them. This entanglement is illustrated by an increase in the logarithmic negativity during the application of spin squeezing. During free evolution, this entanglement decreases due to environmental noise. Thus, the creation of spin squeezing in our spin-1 systems is accompanied by the creation of entanglement.

### v2

@@
 It has been extensively studied in qubit systems [23, 24, 25, 26] and has also been discussed for qutrit systems [27, 28].
-In this study, we explore the potential enhancements in quantum metrological performance achievable by transitioning from two-level (qubits) to three-level (qutrits) spin systems and incorporating spin squeezing within typical quantum magnetometry settings, specifically NV center Ramsey interferometry.
+In this article, we explore the potential enhancements in quantum metrological performance achievable by transitioning from two-level (qubits) to three-level (qutrits) spin systems and incorporating spin squeezing within typical quantum magnetometry settings, specifically NV center Ramsey interferometry.
 We adopt a foundational approach, comparing qubit versus qutrit Ramsey quantum magnetometry scenarios, both with and without spin squeezing, in terms of Quantum Fisher Information (QFI).
@@
 The NV center, characterized as an optically active color defect, results from a substitutional nitrogen impurity and a neighboring carbon vacancy in the diamond lattice.
-Comprising six electrons from nitrogen and surrounding carbon atoms, the negatively charged NV center is described as a spin-1 system.
+Comprising six electrons from nitrogen and the surrounding carbon atoms, the negatively charged NV center is described as a spin-1 system.
 We opt to explore the full capabilities of the NV center by modeling it as a qutrit system.
@@
 1.
-Figure 1: Illustration of Ramsey-type sequence with and without introduction of spin squeezing. Note that the times mentioned in the figure correspond to those considered in our study.
+Figure 1: The application of Ramsey-type sequence on two Nitrogen-Vacancy (NV) Centers with and without the introduction of spin squeezing.
+$T_{\mathrm{fr}}$ is the free evolution time.
+$T_{\mathrm{sq}}$ is the duration of the application of spin squeezing.
 II.1 Free Hamiltonian and Initial states
@@
 Then, each NV center undergoes a $\pi/2$ pulse, achieved by applying a rotation operator,
-$\mathrm{e}^{-i\frac{\pi}{2}S_{x}}.$
+$\mathrm{e}^{-i\frac{\pi}{2}\hat{S}_{x}}.$
 This pulse is experimentally generated by microwave drives [35].
-When incorporating spin squeezing (details provided in the following subsections), the initial state, |$\ket{\psi_{0}}$, is first driven through spin squeezing, and then subjected to the $\pi/2$ pulse, as illustrated in Fig.
-1.
+When incorporating spin squeezing, the initial state, |$\ket{\psi_{0}}$, is first driven through spin squeezing, and then subjected to the $\pi/2$ pulse, as illustrated in Fig. 1.
 II.2 Spin Squeezing Generation
@@
 The choice of $x$-axis twisting in the context of spin squeezing is strategic, as it allows for the manipulation of the noise cone, effectively narrowing the error ellipse of the spin.
-This, combined with the optimal ground state, enhances the precision of the magnetic field measurement by interplaying with the $z$-axis rotation induced by the magnetic field [37].
+This, combined with the optimal initial state, enhances the precision of the magnetic field measurement by interplaying with the $z$-axis rotation induced by the magnetic field [37].
 The generation of this type of Hamiltonian in real setup has been discussed in the literature.
@@
 The method presents a generic scheme for creating spin squeezing in any non-squeezing systems through the use of periodic rotation pulses.
-Figure 2: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G.
+Figure 2: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G for the spin-1 formalism.
 The curves represents the free evolution as follows: the solid blue curve corresponds to a single NV center, the dashed-dotted red curve corresponds to two NV centers, the pointed cyan curve represents two NV centers subjected to spin squeezing ($c_{1}=2.5$ GHz, $T_{\mathrm{sq}}=3$ ns), and the dashed green curve represents two NV centers subjected to spin squeezing ($c_{1}=0.25$ GHz, $T_{\mathrm{sq}}=30$ ns).
@@
 We assume very low temperatures relative to the NV center transition frequencies.
-Dephasing and thermalization are incorporated through the dissipative rate.
-They are respectively given as follows,
+Dephasing and thermalization are incorporated through the dissipators given as follows,
 $\mathcal{D}_{d}(\rho)$ $=\sum^{N}_{i=1}\gamma_{\mathrm{d}}\,[\hat{S}_{zi}\rho(t)\hat{S}_{zi}-\rho(t)];$ (8)
@@
 $\hat{J}_{z}$ $=\frac{1}{2}(\hat{\sigma}_{z}\otimes I_{3}+I_{3}\otimes\hat{\sigma}_{z}),$ (20)
-with $I_{3}$ being the three dimensional unit matrix.
-To calculate spin squeezing, we calculate the mean spin squeezing direction $\hat{n}$ for the state and then take the minimal variance in the perpendicular direction of $\hat{n}$,
+with $I_{3}$ being the three-dimensional unit matrix. To calculate spin squeezing, we calculate the mean spin squeezing direction $\hat{n}$ for the state,
 $\hat{n}=\frac{(\langle\hat{J}_{x}\rangle,\langle\hat{J}_{y}\rangle,\langle\hat{J}_{z}\rangle)}{|\langle\hat{J}\rangle|}.$ (21)
+And then, we take the minimal variance, $(\Delta\hat{J}^{2}_{n_{\perp}})_{\text{min}}$, in the perpendicular direction of $\hat{n}$.
 The presence of spin squeezing is characterized by $\xi^{2}<1$.
@@
 Our results are consistent with this statement (Appendix A).
-Figure 3: The curve in solid blue represents the squeezing parameter, $\xi^{2}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
+Figure 3: For the spin-1 formalism, the curve in solid blue represents the squeezing parameter, $\xi^{2}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
 IV Results
-Initially, we investigate the typical Ramsey sequence for one and two NV centers, both starting in the initial state (Eq.
-5).
-Subsequently, after a $\pi/2$ pulse, we allow the system to undergo evolution through the free Hamiltonian $\hat{H}_{0}$ for a duration of $T_{\mathrm{fr}}=20$ $\mu$s, $T_{\mathrm{fr}}$ being the free evolution time.
-We choose this time taking into account the dissipative time that falls within the order of tenth of microseconds.
-To further enhance precision, starting in the initial state (Eq.
-5), we consider introducing spin-squeezing before the $\pi/2$ pulse and the free evolution.
-The two scenarios are illustrated in Fig.
-1.
+[…]
 IV.1 Analysis with spin-1 formalism
-[…]
 Fig.
-2 compares two NV centers under different conditions: one without spin squeezing and the other with spin squeezing applied at varying strengths, specifically with squeezing constants of $c_{1}=0.25$ GHz and $c_{1}=2.5$ GHz.
-In the absence of noise, the QFI is larger when spin squeezing is applied, and it increases further with a higher squeezing constant, $c_{1}$ (Fig.
-2a).
-When environmental noise is introduced with $\gamma_{\text{d}}=0.1\gamma_{\text{t}}$, spin squeezing remains beneficial, particularly at a lower squeezing constant ( $c_{1}=0.25$ GHz) (Fig.
-2b).
-Applying moderate squeezing, sufficiently large to achieve its effects, shows advantages by being more robust against the impact of environmental noise.
-However, within the first 4 $\mu$s, a higher squeezing constant provides more benefit.
-As the dephasing rate increases to $\gamma_{\text{d}}=\gamma_{\text{t}}$, the advantage of squeezing is lost (Fig.
-2c).
+2 compares one NV center without spin squeezing (blue), two NV centers without spin squeezing (red) and two NV centers with spin squeezing applied at varying strengths, specifically with squeezing constants of $c_{1}=0.25$ GHz (green) and $c_{1}=2.5$ GHz (cyan).
+In the case without spin squeezing, Fig.
+2a shows a twofold increase in the QFI when transitioning from one NV center to two NV centers.
+When environmental noise is introduced (Fig.
+2b and Fig.
+2c), whether the QFI increases or decreases for one NV center, the twofold enhancement persists for two NV centers.
+In the case with spin squeezing, in the absence of noise (Fig.
+2a), the QFI is larger when spin squeezing is applied with squeezing constant, $c_{1}=0.25$ GHz.
+It increases further with a higher squeezing constant, $c_{1}=2.5$ GHz.
+When environmental noise is introduced with $\gamma_{\text{d}}=0.1\gamma_{\text{t}}$ (Fig.
+2b), spin squeezing remains beneficial, particularly at a lower squeezing constant, $c_{1}=0.25$ GHz.
+This suggests that in the presence of noise, a higher squeezing constant does not necessarily lead to higher QFI.
+Instead, a low squeezing constant ($c_{1}=0.25$ GHz), which is sufficient enough to generate spin squeezing, is more advantageous.
+However, within the first 4 $\mu$s, a higher squeezing constant, $c_{1}=2.5$ GHz, provides more benefit.
+Finally, Fig.
+2c shows that the advantage of squeezing is lost when the dephasing rate increases to $\gamma_{\text{d}}=\gamma_{\text{t}}$.
+While spin squeezing enhance magnetic sensing capabilities, they are also highly sensitive to environmental noise, particularly dephasing.
+Analyzing the behavior of the QFI curves under such noise conditions is crucial.
+By identifying the time at which the QFI reaches its maximum, we can optimize experimental design and ensure measurements are performed at the peak sensitivity.
+Experimentally, we typically measure populations or ($\hat{S}_{z}$) values.
+However, it is important to note that while the QFI is calculated for the optimal Positive Operator-Valued Measure (POVM), specifically the eigenprojectors of the SLD, measuring $\hat{S}_{z}$ could potentially be suboptimal.
 The effect of spin squeezing on the NV centers is further illustrated in Fig.
@@
 Appendix A demonstrates that spin squeezing also leads to entanglement.
-The duration of the application of spin squeezing, denoted by $T_{\mathrm{sq}}$, corresponds to the time when the squeezing parameter reaches its minimum value.
-The initial state is oriented along the $z$-axis.
-When spin squeezing is applied, it reduces the variance in a specific direction, creating an ellipsoid shape that remains aligned with the $z$-axis.
+Furthermore, given the initial state is oriented along the $z$-axis, when spin squeezing is applied, it reduces the variance in a specific direction, creating an ellipsoid shape that remains aligned with the $z$-axis.
 A $\pi/2$ pulse then shifts the spin into the $x$-$y$ plane while preserving this variance ellipsoid.
@@
 Finally, at $t_{3}$, a $\pi/2$ pulse is applied, shifting the state into the $x$-$y$ plane.
-Experimentally, we typically measure populations or ($\hat{S}_{z}$) values.
-However, it is important to note that while the QFI is calculated for the optimal Positive Operator-Valued Measure (POVM), specifically the eigenprojectors of the Symmetric Logarithmic Derivative (SLD), measuring $\hat{S}_{z}$ could potentially be suboptimal.
-While spin squeezing enhance magnetic sensing capabilities, they are also highly sensitive to environmental noise, particularly dephasing.
-Analyzing the behavior of the QFI curves under such noise conditions is crucial.
-By identifying the time at which the QFI reaches its maximum, we can optimize experimental design and ensure measurements are performed at the peak sensitivity.
-Figure 5: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G.
+Figure 5: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G for the spin-1/2 formalism.
 The curves represents the free evolution as follows: the solid blue curve corresponds to a single NV center, the dashed-dotted red curve corresponds to two NV centers, the pointed cyan curve represents two NV centers subjected to spin squeezing ($c_{1}=2.5$ GHz, $T_{\mathrm{sq}}=3$ ns), and the dashed green curve represents two NV centers subjected to spin squeezing ($c_{1}=0.25$ GHz, $T_{\mathrm{sq}}=30$ ns).
 The different scenarios are illustrated as: (a) without including any noise terms, (b) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=0.1\gamma_{t}$, and (c) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=\gamma_{t}$.
-The NV centers system are described by spin-1/2 formalism.
 IV.2 Analysis with spin-1/2 formalism
@@
 $\mathrm{e}^{-i\frac{\pi}{2}\hat{\sigma}_{x}}.$
-As in the spin-1 formalism, if spin squeezing is added, it is applied before the $\pi/2$ pulse. We then proceed to analyze the QFI under both scenarios, considering the presence and absence of noise.
+As in the spin-1 formalism, if spin squeezing is added, it is applied before the $\pi/2$ pulse.
+We then proceed to analyze the QFI under both scenarios (with and without squeezing), considering the presence and absence of noise.
 In Fig.
@@
 5b), and Ramsey sequence without spin squeezing is more beneficial.
-In the case of the qutrit, we observe advantages even in the presence of dephasing (Fig.
-2b).
-Therefore, we can see the benefits of using qutrit sensors over qubit sensors.
+Earlier in Fig.
+2b, for the qutrit case, we observed advantages even in the presence of dephasing.
+Therefore, the benefits of using qutrit sensors over qubit sensors are clearly evident.
 V Conclusion
@@
 Throughout our investigation, we have analyzed the dynamics of Quantum Fisher Information (QFI) across these scenarios, utilizing the Lindblad master equation to model the spin system’s evolution under dephasing and thermalizing noise.
-Incorporating a spin squeezing step in magnetometric measurement and utilizing qutrit NV centers have demonstrated a precision increase relative to standard Ramsey magnetometry with qubits and without spin squeezing.
-However, the advantage of qutrit spin squeezing is notably sensitive to the dephasing environment.
+Incorporating a spin squeezing step in magnetometric measurement and utilizing qutrit NV centers have demonstrated a precision increase relative to standard Ramsey magnetometry.
+However, the advantage of spin squeezing is notably sensitive to the dephasing environment.
 Our numerical analysis showed that, as dephasing increases, the advantages become marginal compared to standard methods.
 We also observed that moderate squeezing results in higher QFI compared to strong squeezing, which is only beneficial for a brief period of free evolution.
+Furthermore, by conducting the entire analysis using the spin-1/2 formalism and comparing it with the spin-1 formalism, we have demonstrated the advantages of utilizing a qutrit system over a qubit system.
+Figure 6: For the spin-1 formalism, the curve in solid blue represents the logarithmic negativity, $E_{N}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
 Additionally, we have observed that spin squeezing is accompanied by the generation of entanglement, aligning with the general expectation of quantum entanglement as a metrological resource.
@@
 Our findings contribute to the broader exploration of quantum technologies beyond qubits, where qutrits and higher-dimensional qubits (qudits) are increasingly considered, particularly in the context of quantum computing [52, 53].
-An intriguing question persists regarding the scalability of these advantages, especially regarding the utilization of NV center qutrit ensembles or the investigation of higher-level systems (qudits).
+An intriguing question persists regarding the scalability, as outlook in [54], of these advantages, especially regarding the utilization of NV center qutrit ensembles or the investigation of higher-level systems (qudits).
 These avenues pose exciting challenges for future studies, promising further insights.
-Figure 6: The curve in solid blue represents the logarithmic negativity, $E_{N}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
+Acknowledgements
+This work was supported by the Scientific and Technological Research Council of Türkiye (TÜBITAK) under Project Number 123F150. L. G. and Ö. E. M. thank TÜBITAK for their support.
 Appendix A Entanglement Measure

### v3

@@
 II.2 Spin Squeezing Generation
-To describe the intricate dynamics and capture the squeezing effects in the Hamiltonian for two NV centers, a squeezing term is introduced [36],
+To describe the intricate dynamics and capture the squeezing effects in the Hamiltonian for two NV centers, a squeezing term is introduced,
 $\hat{H}_{\mathrm{S}}=c_{1}\left(\sum_{i=1}^{N}\hat{S}_{xi}\right)^{2},$ (6)
@@
 The Hamiltonian, as depicted in Eq.
-(6), is referred as a one-axis twisting Hamiltonian.
+(6), is referred as a one-axis twisting Hamiltonian [20, 36].
 The choice of $x$-axis twisting in the context of spin squeezing is strategic, as it allows for the manipulation of the noise cone, effectively narrowing the error ellipse of the spin.
 This, combined with the optimal initial state, enhances the precision of the magnetic field measurement by interplaying with the $z$-axis rotation induced by the magnetic field [37].
-The generation of this type of Hamiltonian in real setup has been discussed in the literature.
+II.2.1 Pulse Sequence Method
+A more direct approach to generating spin squeezing is through pulse sequences that induce an axis-twisting effect.
+In [21] a method for achieving spin squeezing via such pulse sequences is described in detail.
+While their work focuses on spin-1/2 systems, this approach can be extended to spin-1 systems.
+This is particularly relevant to our setup, where we are interested in generating x-axis twisting.
+By adjusting the pulse properties, the method should allow us to transit from z-axis to x-axis twisting, which is more pertinent to the spin dynamics we are studying.
+A recalculation of the squeezing parameters within our parameter ranges will be necessary, but the method is sufficiently flexible to be implemented within our protocol.
+II.2.2 Electric Field Induced Strain Method
+An alternative method for generating spin squeezing, besides using pulse sequences, is the application of a strong electric field to induce strain in the system, leading to a Hamiltonian of the form $\sim[(S_{x})^{2}-(S_{y})^{2}]$.
+Although this Hamiltonian differs from the one used in our work, it still develops spin squeezing along specific axes, as demonstrated in [38].
+II.2.3 Off-resonant Field Mediated Effective Interaction Method
+Another approach, commonly used in spin-1/2 systems, involves effective interactions of the form $\sim S^{+}S^{-}+h.c.$, mediated by an off-resonant field that is adiabatically eliminated, as outlined in [39].
+To generalize this to spin-1 systems, one can introduce strong laser fields and replace photon operators with c-numbers, yielding a Hamiltonian of the form $\sim g(S^{+})^{2}+h.c.$ Alternatively, large detuning can be employed via the Schrieffer-Wolff (SW) transformation, also known as the Fröhlich transformation, as shown in [40], to eliminate photon degrees of freedom, leading to a simpler effective Hamiltonian.
+This method has also been applied in systems like silicon-vacancy centers, which are similar to NV centers, as described in [41].
+II.2.4 Other Methods
+The generation of this type of Hamiltonian in real setup has been further discussed in the literature.
 It can be realized through phonon mediation, as evidenced in [42] or produced in optical cavities [43].
@@
 A dynamic control method for NV center spin squeezing, as discussed in [44], could potentially address the technical challenge.
-Alternative approaches to generating spin squeezing have been proposed.
-One example involves using microwave fields in silicon-vacancy centers, which are similar to NV centers [41].
-Additionally, a universal method, introduced in [21], is also applicable to NV centers.
-The method presents a generic scheme for creating spin squeezing in any non-squeezing systems through the use of periodic rotation pulses.
+Recent work [45] demonstrates the generation of large-scale spin squeezing from finite-temperature easy-plane magnetism, a method that shows great promise for NV centers.
+To approximate a realistic measurement scenario, we calculate the Classical Fisher Information in Appendix B.
 Figure 2: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G for the spin-1 formalism.
@@
   • during spin squeezing by $\hat{H}=\hat{H}_{0}+\hat{H}_{\mathrm{S}}$.
+To model the system, we adopt a Markovian noise framework and analyze the interplay between dephasing and thermalization coefficients in optimizing the squeezing effect.
+Our approach employs a phenomenological master equation incorporating decoherence and dephasing terms, consistent with standard treatments of bulk NV centers, which rely on jump operators linear in spin-1 defect operators ($S_{+}$, $S_{-}$, and $S_{z}$) [47, 48, 49].
 The NV center, while a robust quantum system, is inevitably influenced by the fluctuating states of nearby nuclear spins.
@@
 […]
+In this study, we focus on NV centers located deep within the bulk of the diamond, at depths exceeding tens of nanometers from the crystal surface.
+This assumption minimizes the influence of surface effects, such as charge noise caused by near-surface defects, allowing us to concentrate on bulk NV center dynamics.
+While charge noise is generally more pronounced for near-surface NV centers, its contribution decreases significantly with increasing depth, as demonstrated in [51].
+This reference provides a detailed quantitative analysis of the depth-dependent impact of charge and spin noise on decoherence and relaxation rates.
+For instance, it shows that the electric field fluctuations due to surface charges scale inversely with the square of the defect depth.
+At depths greater than tens of nanometers, the noise contribution from surface charges becomes negligible compared to bulk noise sources.
+These findings support our assumption that the charge noise for deep NV centers is substantially less impactful.
+Additionally, experimental results in [52] confirm the depth dependence of relaxation and dephasing times, illustrating a marked improvement in coherence times as NV center depth increases.
+Figures 1 and 2 of [52] highlight this trend, with T2 and T1 coherence times improving by approximately two orders of magnitude as the depth increases from 4 nm to 12 nm.
 III Quantum Measures
@@
 where $n_{\perp}$ is an axis perpendicular to the mean-spin direction $\hat{n}$, $N$ is the number of spins (in our case, $N=2$), and $\hat{J}$ represents the collective spin operator.
+In this work, we are using the particle-entanglement-based definition of spin squeezing, which is more restrictive than the noise redistribution version introduced by Wineland [56], commonly used in Ramsey measurements.
+The entanglement-based definition requires true multi-particle entanglement and cannot be achieved with a single spin, whereas the Wineland definition can reflect spin squeezing even in the absence of particle entanglement, merely by redistributing noise between different spin components.
 Kitagawa-Ueda and other quantum correlations based criteria of spin squeezing [57] are commonly defined for SU(2) or spin-1/2 atomic ensembles.
@@
 Our results are consistent with this statement (Appendix A).
+Our approach highlights the quantum metrological advantage of employing higher-dimensional (spin-1) multipartite quantum entanglement.
+This advantage, distinct from conventional spin noise squeezing, is achieved by leveraging entanglement-based criteria for spin squeezing to further enhance sensitivity in quantum measurement protocols.
 Figure 3: For the spin-1 formalism, the curve in solid blue represents the squeezing parameter, $\xi^{2}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
@@
 However, it is important to note that while the QFI is calculated for the optimal Positive Operator-Valued Measure (POVM), specifically the eigenprojectors of the SLD, measuring $\hat{S}_{z}$ could potentially be suboptimal.
+Our method is designed for measuring DC magnetic fields, where the typical readout technique is based on photoluminescence [26].
+Additionally, optically detected magnetic resonance (ODMR) is commonly employed in such measurements [27, 61, 62].
 The effect of spin squeezing on the NV centers is further illustrated in Fig.
@@
 Appendix A demonstrates that spin squeezing also leads to entanglement.
-Furthermore, given the initial state is oriented along the $z$-axis, when spin squeezing is applied, it reduces the variance in a specific direction, creating an ellipsoid shape that remains aligned with the $z$-axis.
-A $\pi/2$ pulse then shifts the spin into the $x$-$y$ plane while preserving this variance ellipsoid.
-The effect of the magnetic field $B$ is to precess the spin around the $z$-axis, and the $x$-axis twisting deforms the noise cone, effectively narrowing the spin’s error ellipse.
-This sequence is illustrated in Fig.
-4.
+The spin system is initially oriented along the $z$-axis. The expectation values of the angular momentum operators are
+$\langle J_{z}\rangle=\frac{N}{2},\quad\langle J_{x}\rangle=0,\quad\langle J_{y}\rangle=0,$ (22)
+where $N$ is the total number of spins in the ensemble.
+In this state, the spin variance is symmetric around the $z$-axis, forming a circular uncertainty distribution in the $x$-$y$-plane.
+The uncertainty in each component is
+$\langle(\Delta J_{x})^{2}\rangle=\langle(\Delta J_{y})^{2}\rangle=\frac{N}{4}.$ (23)
+When spin squeezing is applied, the variance along one axis in the $x$-$y$ plane is reduced, while the variance along the orthogonal axis increases. The squeezed variances become,
+$\langle(\Delta J_{\text{sq}})^{2}\rangle=\frac{N}{4\xi^{2}},\quad\langle(\Delta J_{\text{anti-sq}})^{2}\rangle=\frac{N\xi^{2}}{4},$ (24)
+where $\xi^{2}<1$ is the squeezing parameter.
+The squeezed and anti-squeezed axes lie in the $x$-$y$ plane, but their exact orientation within the plane may not be specified.
+After applying a $\pi/2$ pulse, the spin state is rotated into the $x$-$y$ plane, and the noise ellipse is now aligned in this plane.
+The pulse does not alter the relative squeezing, but just the orientation of the state.
+The effect of the magnetic field $B$, applied along the $z$-axis, is to precess the spin state around the $z$-axis.
+While we did not adopt a non-exponential decoherence model, such an approach could provide valuable insights by capturing slower, more realistic noise dynamics [63].
+Furthermore, non-exponential decay could potentially extend coherence times, thus enhancing metrological precision in practical applications.
+Investigating this further may increase the relevance of our results to real-world quantum systems.
+Moreover, we examine non-Markovian effects by introducing an ancillary qubit into the model, resulting in improved metrological performance, as discussed in Section IV.3.
 Figure 4: Illustration of the mean spin orientation and the variance of state preparation before the free evolution of the NV centers.
@@
 Therefore, the benefits of using qutrit sensors over qubit sensors are clearly evident.
+IV.3 Non-Markovian Environment
+To investigate non-Markovian effects in our system [64, 65], we introduce an ancillary qubit as a proof of principle.
+In our setup, this qubit is coupled to a single NV center through an $XX+YY$ interaction with a coupling strength of 0.1 GHz.
+The ancillary qubit, much like the NV center, is influenced by environmental effects such as dephasing and thermalization.
+However, in this case, the NV center experiences these environmental influences indirectly through its interaction with the ancillary spin.
+By adding this single ancillary qubit, we effectively create a finite spin environment for the NV center.
+This finite environment can mimic a structured, non-Markovian bath, shifting the system’s response from a typical Markovian to a non-Markovian regime.
+As shown in Fig.
+6, the QFI is significantly higher in the non-Markovian case, demonstrating enhanced sensitivity and suggesting that such structured environments may be advantageous for metrological applications.
+In practice, NV centers are often surrounded by a finite number of nuclear or impurity spins, which can act as a non-Markovian environment and offer similar metrological benefits.
+By introducing an ancillary qubit as an effective finite environment, we provide a simplified model that captures these effects, highlighting the potential of non-Markovian dynamics for enhanced quantum sensing [66, 67].
+Figure 6: For the spin-1 formalism, the red dotted curve shows the quantum Fisher information (QFI) as a function of time during the free evolution of a single NV center, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
+The blue solid curve represents QFI under Markovian noise, while the red solid curve shows QFI under non-Markovian noise.
 V Conclusion
@@
 Figure 7: For the spin-1 formalism, the curve in solid blue represents the logarithmic negativity, $E_{N}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
+In light of surface effects, the dominant role of charge noise near the surface severely challenges the exploitation of spin squeezing and multipartite entanglement for magnetometry.
+Addressing these limitations requires exploring alternative strategies to enable quantum entanglement advantages for near-surface probes.
+Future research could focus on identifying optimal depths and surface properties to enhance magnetometry precision under non-Markovian dynamics and spin squeezing.
+These efforts are crucial for unlocking the full potential of NV centers for advanced sensing applications.
 Additionally, we have observed that spin squeezing is accompanied by the generation of entanglement, aligning with the general expectation of quantum entanglement as a metrological resource.
@@
 In Fig.
-6, we observe that the application of spin squeezing to the two NV centers system creates an entanglement between them.
+7, we observe that the application of spin squeezing to the two NV centers system creates an entanglement between them.
 This entanglement is illustrated by an increase in the logarithmic negativity during the application of spin squeezing.
@@
 Thus, the creation of spin squeezing in our spin-1 systems is accompanied by the creation of entanglement.
+Appendix B Classical Fisher Information
+We calculate the Classical Fisher Information (CFI) during the system’s free evolution to better reflect a realistic experimental setup.
+Specifically, we evaluate the CFI using measurements along the eigenvectors of $S_{x}$ and $S_{y}$, which can be practically implemented through state tomography to reconstruct the quantum state and analyze its sensitivity to parameter changes.
+In quantum metrology, the CFI is a valuable metric for quantifying how responsive a measurement is to variations in a system parameter.
+For a probability distribution $\{p_{i}(\theta)\}$ that depends on a parameter $\theta$, the CFI is given by
+Figure 8: The dark blue dashed line represents the Quantum Fisher Information (QFI) for a two-NV-center spin-1 system under spin squeezing conditions, with parameters $c_{1}=0.25$ GHz and $T_{\text{sq}}=30$ ns, accounting for thermalization and dephasing rates of $\gamma_{t}=0.2$ MHz and $\gamma_{d}=0.1\,\gamma_{t}$.
+The red dashed line indicates the QFI without spin squeezing.
+The cyan dotted line shows the Classical Fisher Information (CFI) using eigenvectors $S_{x}$ and $S_{y}$ as the POVM for a two-NV-center spin-1 system with spin squeezing (parameters: $c_{1}=0.25$ GHz, $T_{\text{sq}}=30$ ns), while the solid magenta line represents the CFI without spin squeezing.
+$\text{CFI}(\theta)=\sum_{i}\frac{1}{p_{i}(\theta)}\left(\frac{\partial p_{i}(\theta)}{\partial\theta}\right)^{2},$ (31)
+where $p_{i}(\theta)$ represents the probability of the $i$-th measurement outcome for a given $\theta$.
+In our context, these probabilities correspond to outcomes measured along the eigenvectors of $S_{x}$ and $S_{y}$, providing the necessary basis for CFI calculation.
+By choosing eigenvectors of $S_{x}$ and $S_{y}$ as the Positive Operator-Valued Measure (POVM) elements, we model measurements that are feasible in practical setups.
+Such measurements can be carried out using state tomography, as detailed in [72].
+In Fig.
+8, we observe that the CFI is non-zero and is higher with squeezing, which further supports our results.